Exercise 15-4 (Part Level Submission)

Bridgeport, Inc. is trying to establish the standard labor cost of a typical oil change. The following data have been collected from time and motion studies conducted over the past month.

Actual time spent on the oil change 1 hour
Hourly wage rate $10
Payroll taxes 10% of wage rate
Setup and downtime 16% of actual labor time
Cleanup and rest periods 34% of actual labor time
Fringe benefits 25% of wage rate

If a oil change took 1.70 hours at the standard hourly rate, what was the direct labor quantity variance? (Round answer to 2 decimal places, e.g. 1.25.)

Direct labor quantity variance $

FavorableNeither favorable nor unfavorableUnfavorable
Answer #1

Actual hours of input used = 1.70

Standard hours = 1 + (1 x 16%) + (1 x 34%) = 1 + 0.16 + 0.34 = 1.50

Standard rate per hour = 10 + (10 x 10%) + (10 x 25%)

= 10 + 1 + 2.5 = 13.50

Direct Labor quantity variance:-

(Actual hours - standard hours) x standard price

= (1.70 - 1.50) x 13.50

= $2.70 unfavorable